OpenStudy (anonymous):

What is a quartic function with only the two real zeros given? x = -4 and x = -1

OpenStudy (anonymous):

x = -4 x + 4 = 0 x = -1 x + 1=0 (x+4)(x+1)=0 x^2+5x+4=0